The residue at z=1 of the function 2z/(1-z)(z+1)


1
Expert's answer
2021-08-12T07:28:16-0400

The residue at z= 1 is

"\lim _{z\to \:1}\left(\left(z-1\right)\frac{2z}{\left(1-z\right)\left(z+1\right)}\right)\\
=-\frac{2z}{z+1}\\
=\lim _{z\to \:1}\left(-\frac{2z}{z+1}\right)\\
\mathrm{Plug\:in\:the\:value}\:z=1\\
=-\frac{2\cdot \:1}{1+1}\\
=-1"